# Dual-Unitary Classical Shadow Tomography

### v1

@@ Abstract
We study operator spreading in random dual-unitary circuits within the context of classical shadow tomography. Primarily, we analyze the dynamics of the Pauli weight in one-dimensional qubit systems evolved by random two-local dual-unitary gates arranged in a brick-wall structure, ending with a final measurement layer. We do this by deriving general constraints on the Pauli weight transfer matrix and specializing to the case of dual-unitarity. We first show that dual-unitaries must have a minimal amount of entropy production. Remarkably, we find that operator spreading in these circuits has a rich structure resembling that of relativistic quantum field theories, with massless chiral excitations that can decay or fuse into each other, which we call left- or right-movers. We develop a mean-field description of the Pauli weight in terms of $\rho(x,t)$, which represents the probability of having nontrivial support at site $x$ and depth $t$ starting from a fixed weight distribution. We develop an equation of state for $\rho(x,t)$, and simulate it numerically using Monte Carlo simulations. Lastly, we demonstrate that the fast-thermalizing properties of dual-unitary circuits make them better at predicting large operators than shallow brick-wall Clifford circuits.
pacs:
Valid PACS appear here
@@ I Introduction
The dynamics of quantum information in closed, thermalizing systems is a core to our understanding of quantum many-body physics Deutsch (1991); Srednicki (1994); Rigol et al. (2008). In thermalizing quantum systems, local quantum information is scrambled between constituent components and the entanglement exhibits volume-law scaling. Initially localized operators spread ballistically and their evolution across space and time can be used to define various measures of information propagation, such as the out-of-time-order (OTOC) correlator Larkin and Ovchinnikov (1969); Lieb and Robinson (1972); Dóra and Moessner (2017); Nahum et al. (2017, 2018); von Keyserlingk et al. (2018); Brown and Fawzi (2013); Styliaris et al. (2021); Zanardi and Anand (2021). The physics of these systems is often contrasted with the phenomena of many-body localization, where a closed, interacting many-body quantum system can evade thermalization, and local quantum information can escape global scrambling Nandkishore and Huse (2015). Rather that studying localization, however, our focus in this paper is restricted to thermalizing systems generated by dual-unitary circuits. Dual-unitary circuits are fast-scrambling circuits that produce chaotic dynamics, and furthermore have a symmetry that enables efficient calculation of observables Borsi and Pozsgay (2022).
@@
Our primary object of study is the Pauli weight of an operator. The Pauli weight of a fixed operator $O$ characterizes how much of it is distributed over different Pauli strings: $w_{O}(P)=\mathop{\operatorname{Tr}}(OP)^{2}$, where $P=\bigotimes_{i}P_{i}$ is a Pauli string i.e. a product of local Pauli operators $P_{i}\in\{\mathds{1},X,Y,Z\}.$ By studying the evolution of the Pauli weight over time and over different Pauli strings, we can analyze how quantum information spreads over space and time, similar to previous works on the operator spreading in random unitary circuits Ho and Abanin (2017); Bohrdt et al. (2017); Nahum et al. (2017); Kukuljan et al. (2017); Nahum et al. (2018); von Keyserlingk et al. (2018); Khemani et al. (2018); Rakovszky et al. (2018); Chan et al. (2018); Zhou and Nahum (2019); Zhou and Chen (2019); Qi et al. (2019); Xu and Swingle (2019); Chen and Zhou (2019); Parker et al. (2019); Kuo et al. (2020); Akhtar and You (2020); von Keyserlingk et al. (2022); Schuster and Yao (2022). The Pauli weight is also interesting from the perspective of classical shadow tomography Huang et al. (2020); Ohliger et al. (2013); Guta et al. (2020); Huang et al. (2021); Hadfield et al. (2022); Elben et al. (2020); Enshan Koh and Grewal (2022); Hu and You (2022); Hu et al. (2023); Levy et al. (2021); Bu et al. (2022); Hu et al. (2022); Seif et al. (2022); Hao Low (2022); Akhtar et al. (2023); Bertoni et al. (2022); Arienzo et al. (2022); Ippoliti et al. (2023); Hu et al. (2024), where it may be used to characterize the sample complexity of predicting the expectation value of a Pauli string $P$ Hu et al. (2023); Akhtar et al. (2023, 2023). In this work, we will use our observations about the Pauli weight in dual-unitary circuits to analyze their usefulness for predicting extensive properties of quantum states.
@@
The primary scientific question we aim to address is how locally-scrambled random dual-unitary circuits differ from conventional random Haar brickwall circuits in their Pauli weight dynamics, and what this implies about the advantages or disadvantages of using dual-unitaries for prediction. By thinking of the Pauli weight as a state vector in a many-body quantum system, we are able to interpret the dynamics in the familiar language of condensed matter physics. We learn that dual-unitary Pauli weight dynamics are unique in that they allow for the existence of stable, chiral excitations which we call left and right-movers, in analogy with the components of the Dirac field. These chiral excitations are freely propagating and are responsible for the ballistic front in the average weight. The chiral movers form the left and right branches of a future light cone when studying the Pauli weight of contiguous operators evolved by random dual-unitary circuits. These chiral movers, and therefore the front of the average Pauli weight, propagate at the “speed of light” ($v=1$) and contribute to the fast-scrambling of quantum information in the system (see for e.g., Refs. Bertini et al. (2020); Holden-Dye et al. (2023) which also discuss chiral excitations in dual-unitary circuits). The rapid rate of thermalization in random dual-unitary circuits makes them superior for predicting extensive operators compared to brick-wall arrangements of Clifford gates.
@@
This is established by evaluating the scaling base of the shadow norm, $\beta$, for extensive operators in dual-unitary brick-wall circuits compared to Clifford brick-wall circuits, where $\beta=(\|P\|^{2})^{1/n}$, and $P$ is an Pauli string operator with support on all $n$ sites. We verify that this $\beta$ can be lower in the dual-unitary case for sufficiently large scrambling parameter $\alpha$ using both numerics and analytical arguments based off of a mean-field approximation. The argument starts by defining a simple order parameter in the statistical model of chiral movers described above: the average occupation $\rho(x,t)$ at each point in space-time $(x,t)$. The average occupation characterizes the probability of finding a particle at a given space-time point. We analyze the dynamics of this order parameter via a mean-field approximation. With our simplified model of the dynamics, it’s clear that the shadow norm of contiguous operators has contributions from two competing effects in the average occupation: saturation, which drives $\rho(x,t)$ from 1 to its equilibrium value of 3/4 and thereby reduces $\beta$; and dispersion, which expands the spatial region where $\rho(x,t)$ is non-zero and thereby increases the shadow norm. Dual-unitaries are able to saturate quickly, but will generally have higher shadow norm for local operators because of their fast dispersion. We are able to eliminate the effect of dispersion by considering extensive operators, which have full or nearly full support, and achieve a lower total shadow norm compared to brick-wall Clifford circuits.
@@
We begin by giving a brief overview of classical shadow tomography as it is relevant to this paper. Next, we introduce the concept of dual-unitarity in circuits and our notion of random, locally-scrambled dual-unitary gates. Equipped with this knowledge, we motivate the set up and the Pauli weight transfer matrix which defines how the Pauli weight evolves as a state vector. Using the definitions of dual-unitary and some information-theoretic bounds, we derive the general form of the transfer matrix in the dual-unitary case. Then, we study the dynamics implied by this transfer matrix as a set of Feynman rules on the left and right-movers. Essentially, the movers are allowed to fuse and by emitted by one another, but one cannot arbitrarily transform into the other. This is the key difference with transfer matrices which are derived form two-designs. We then introduce the average occupation $\rho(x,t)$, which characterizes the average Pauli weight on a single site, and we study its structure via Monte Carlo simulations. Then, we derive a mean-field equation of state of the occupation, argue for why this description remains accurate, and analyze the equation of state and solve it in the homogeneous case. Lastly, we argue how these dual-unitary brick-wall circuits may be advantageous compared to Clifford brick-wall circuits for the prediction of extensive observables using classical shadow tomography.
@@ II.1.1 General Formalism
Figure 1: (a). The traditional set up for classical shadow tomography. To perform prediction on the quantum state $\rho$, we evolve by random unitary $U$ chosen from the ensemble $\mathcal{E}_{U}$. At the end, we perform a final layer of measurements and store the outcomes. In post-processing, we use the classical shadows $\hat{\rho}$ as least-squares estimators of the state $\rho$ evaluated on certain observables. (b). The circuit which evaluates the Pauli weight $w_{\mathcal{E}_{\sigma}}(P)$ given the support of the Pauli string $P$. The middle section of brick-wall transfer matrices $T$ (purple) matches the structure of $\mathcal{E}_{U}$. Each transfer matrix $T$ defines a local Markov process, so that the resulting “state” before the final measurement defines a probability distribution over possible supports of the Pauli string. See Eq. (27) for the general form of the transfer matrix. The final layer of measurement gates $T_{M}$ (red) projects the resulting Pauli weight distribution onto ${|0\rangle}+\frac{1}{3}{|1\rangle}$ where-ever a measurement was performed on the quantum state.
@@
Traditionally, the interest in classical shadow tomography has been mostly geared towards experimental applications: predicting linear observables of generic states on quantum devices. However, in this work, we wish to use the general set up of classical shadow tomography as a backdrop to studying operator spreading in dual-unitary circuits. There are two primary reasons we take this approach: one, this scheme is quite general and allows us to adjust the circuit depth $t$, the system size $n$, the type of gate applied, etc.; two, the growth of the Pauli weight, which characterizes the spatial distribution of an operator, is closely related to the measurement channel of the classical shadow scheme; third, operator spreading in this context characterizes the sample complexity of predicting said operator with respect to the particular circuit model via the shadow norm. Hence, in this section we will introduce the relevant concepts of classical shadow tomography.
@@
Classical shadow tomography was introduced as an efficient protocol for predicting quantum state observables based on few measurements Huang et al. (2020). The scheme works by generating efficient classical representations (or “snapshots”) of the original state, usually stabilizer states, by repeatedly applying a random unitary to the original state and measuring all the qubits afterwards. The specific random unitary ensemble which is used (e.g. ${\mathcal{C}\ell}(2)^{\otimes n},{\mathcal{C}\ell}(2^{n}),$ etc.) may be suited for certain types of observables over others.
@@
The protocol consists of two steps: randomized measurement and classical post processing, as illustrated in Fig. 1(a). Starting with the initial state $\rho$, apply a randomly sampled unitary $U\in\mathcal{E}_{U}$ so that $\rho\rightarrow U\rho U^{\dagger}$. Then perform a projective measurement on the transformed state $U\rho U^{\dagger}$ in the computational basis ($Z$-basis on each qubit independently). The resulting bit-string state ${|b\rangle}$ (labeled by the bit-string $b$ of measurement outcomes), as well as the performed unitary transformation $U$, will be recorded. These two pieces of information define a snapshot state $\hat{\sigma}=U^{\dagger}{|b\rangle}{\langle b|}U$. The randomized measurement protocol can be formulated as a quantum channel $\mathcal{M}$, which maps the initial state $\rho$ to the average snapshot state $\sigma$ by
@@
If the unitary ensemble $\mathcal{E}_{U}$ is local-basis invariant (or equivalently, Pauli-invariant), then the eigen-operators of the measurement channel $\mathcal{M}$ are Pauli strings $P$, with eigenvalue $w_{\mathcal{E}_{\sigma}}$ denoted as the Pauli weight Bu et al. (2022); Hu et al. (2023); Akhtar et al. (2023, 2023). Furthermore, this Pauli weight is not dependent on the specific Pauli string but on its support. It is named as such because it characterizes the “weight” of an operator on a particular sub-region defined by the support of the Pauli,
@@
is the locally-scrambled shadow norm $\|P\|_{\mathcal{E}_{\sigma}}^{2}$, which characterizes the sample complexity of predicting $P$ using the tomography scheme defined by $\mathcal{E}_{\sigma}$ Hu et al. (2023); Akhtar et al. (2023).
@@ II.1.2 Dual-Unitarity in Qubits
We are interested in simulating the evolution of an initial Pauli string $P$ under a brick-wall configuration of random, dual-unitary gates as depicted in Fig. 1. To define the ensemble average and second moment of random, dual-unitary gates, we must first define what a dual-unitary gate is. Dual-unitary circuits have generated a lot of buzz because they’re local; they’re fast scramblers; they can generate chaotic dynamics; they can be used to form $k$-designs; and, they have simple, tractable expressions for two- and four-point correlation functions, including out-of-time-ordered correlators (OTOCs) Borsi and Pozsgay (2022).
@@
Consider a $n$-qubit system arranged on a lattice with Hilbert space $\mathcal{H}=\otimes_{i}\mathcal{H}_{i}$, where $\dim\mathcal{H}_{i}=2$. Every two-local dual-unitary qubit gate $U$ can be parameterized (up to an overall phase) by four arbitrary single-qubit unitaries $u_{\pm},v_{\pm}$, and a single coefficient $J$, which tunes the coupling in a $XXZ$-type interaction Hamiltonian Bertini et al. (2019a, b); Piroli et al. (2020); Claeys and Lamacraft (2021); Kos and Styliaris (2023). This is given below,
@@
We are interested in the average, basis independent behavior of dual-unitary circuits for prediction in a classical shadow setting. If we choose $u_{\pm},v_{\pm}$ according to a random unitary ensemble, then we are able to define a random, dual-unitary gate ensemble $\mathcal{E}_{U}$.

### v2

@@ Abstract
We introduce a classical shadow tomography scheme based on dual-unitary brick-wall circuits termed “dual-unitary shadow tomography” (DUST). For this we study operator spreading and Pauli weight dynamics in one-dimensional qubit systems, evolved by random two-local dual-unitary gates arranged in a brick-wall structure, ending with a final measurement layer. We do this by deriving general constraints on the Pauli weight transfer matrix and specializing to the case of dual-unitarity. We first show that dual-unitaries must have a minimal amount of entropy production. Remarkably, we find that operator spreading in these circuits have a rich structure resembling that of relativistic quantum field theories, with massless chiral excitations that can decay or fuse into each other, which we call left- or right-movers. We develop a mean-field description of the Pauli weight in terms of $\rho(x,t)$, which represents the probability of having nontrivial support at site $x$ and depth $t$ starting from a fixed weight distribution. We develop an equation of state for $\rho(x,t)$, and simulate it numerically using Monte Carlo simulations. Lastly, we demonstrate that the fast-thermalizing properties of dual-unitary circuits make them better at predicting large operators than shallow brick-wall Clifford circuits. Our results are robust to finite-size effects due to the chirality of dual-unitary brick-wall circuits.
@@ I Introduction
The dynamics of quantum information in closed, thermalizing systems is a core to our understanding of quantum many-body physics [1, 2, 3]. In thermalizing quantum systems, local quantum information is scrambled between constituent components and the entanglement exhibits volume-law scaling. Initially localized operators spread ballistically and their evolution across space and time can be used to define various measures of information propagation, such as the out-of-time-order (OTOC) correlator [4, 5, 6, 7, 8, 9, 10, 11, 12]. The physics of these systems is often contrasted with the phenomena of many-body localization, where a closed, interacting many-body quantum system can evade thermalization, and local quantum information can escape global scrambling [13]. Rather than studying localization, however, our focus in this paper is restricted to thermalizing systems generated by dual-unitary circuits. Dual-unitary circuits are fast-scrambling circuits that produce chaotic dynamics and furthermore have a symmetry that enables efficient calculation of observables [14].
@@
Our primary object of study is the Pauli weight of an operator. The Pauli weight of a fixed operator $O$ characterizes how much of it is distributed over different Pauli strings: $w_{O}(P)=\mathop{\operatorname{Tr}}(OP)^{2}$, where $P=\bigotimes_{i}P_{i}$ is a Pauli string i.e. a product of local Pauli operators $P_{i}\in\{\mathds{1},X,Y,Z\}.$ By studying the evolution of the Pauli weight over time and over different Pauli strings, we can analyze how quantum information spreads over space and time, similar to previous works on the operator spreading in random unitary circuits [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. The Pauli weight is also interesting from the perspective of classical shadow tomography [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52], where it may be used to characterize the sample complexity of predicting the expectation value of a Pauli string $P$ [42, 48, 53]. In this work, we will use our observations about the Pauli weight in dual-unitary circuits to analyze their usefulness for predicting extensive properties of quantum states.
@@
The primary scientific question we aim to address is how locally-scrambled random dual-unitary circuits differ from conventional random Haar brick-wall circuits in their Pauli weight dynamics, and what this implies about the advantages or disadvantages of using dual-unitaries for prediction. By thinking of the Pauli weight as a state vector in a many-body quantum system, we are able to interpret the dynamics in the familiar language of condensed matter physics. We learn that dual-unitary Pauli weight dynamics are unique in that they allow for the existence of stable, chiral excitations which we call left and right-movers, in analogy with the components of the Dirac field. These chiral excitations are freely propagating and are responsible for the ballistic front in the average weight. The chiral movers form the left and right branches of a future light cone when studying the Pauli weight of contiguous operators evolved by random dual-unitary circuits. These chiral movers, and therefore the front of the average Pauli weight, propagate at the “speed of light” ($v=1$) and contribute to the fast-scrambling of quantum information in the system (see for e.g., Refs. [54, 55] which also discuss chiral excitations in dual-unitary circuits). The rapid rate of thermalization in random dual-unitary circuits makes them superior for predicting extensive operators compared to brick-wall arrangements of Clifford gates.
@@
This is established by evaluating the scaling base of the shadow norm, $\beta$, for extensive operators in dual-unitary brick-wall circuits compared to Clifford brick-wall circuits, where $\beta=(\|P\|^{2})^{1/n}$, and $P$ is a Pauli string operator with support on all $n$ sites. We verify that this $\beta$ can be lower in the dual-unitary case for sufficiently large scrambling parameter $\alpha$ using both numerics and analytical arguments based off of a mean-field approximation. The argument starts by defining a simple order parameter in the statistical model of chiral movers described above: the average occupation $\rho(x,t)$ at each point in space-time $(x,t)$. The average occupation characterizes the probability of finding a particle at a given space-time point. We analyze the dynamics of this order parameter via a mean-field approximation. With our simplified model of the dynamics, it’s clear that the shadow norm of contiguous operators has contributions from two competing effects in the average occupation: saturation, which drives $\rho(x,t)$ from 1 to its equilibrium value of 3/4 and thereby reduces $\beta$; and dispersion, which expands the spatial region where $\rho(x,t)$ is non-zero and thereby increases the shadow norm. Dual-unitaries are able to saturate quickly, but will generally have higher shadow norm for local operators because of their fast dispersion. We are able to eliminate the effect of dispersion by considering extensive operators, which have full or nearly full support, and achieve a lower total shadow norm compared to brick-wall Clifford circuits.
@@
We begin by giving a brief overview of classical shadow tomography as it is relevant to this paper. Next, we introduce the concept of dual-unitarity in circuits and our notion of random, locally-scrambled dual-unitary gates. Equipped with this knowledge, we motivate the setup and the Pauli weight transfer matrix which defines how the Pauli weight evolves as a state vector. Using the definitions of dual-unitary and some information-theoretic bounds, we derive the general form of the transfer matrix in the dual-unitary case. Then, we study the dynamics implied by this transfer matrix as a set of Feynman rules on the left and right-movers. Essentially, the movers are allowed to fuse and by emitted by one another, but one cannot arbitrarily transform into the other. This is the key difference with transfer matrices which are derived form two-designs. We then introduce the average occupation $\rho(x,t)$, which characterizes the average Pauli weight on a single site, and we study its structure via Monte Carlo simulations. Then, we derive a mean-field equation for the dynamics of the average occupation, argue for why this description remains accurate, analyze the equation of state, and solve it in the homogeneous case. Lastly, we argue how these dual-unitary brick-wall circuits are advantageous compared to Clifford brick-wall circuits for the prediction of extensive observables using classical shadow tomography. This introduces our dual-unitary shadow tomography (DUST) scheme.
@@ II.1.1 General Formalism
Figure 1: (a). The traditional setup for classical shadow tomography. To perform prediction on the quantum state $\rho$, we evolve by random unitary $U$ chosen from the ensemble $\mathcal{E}_{U}$. At the end, we perform a final layer of measurements and store the outcomes. In post-processing, we use the classical shadows $\hat{\rho}$ as least-squares estimators of the state $\rho$ evaluated on certain observables. (b). The circuit that evaluates the Pauli weight $w_{\mathcal{E}_{\sigma}}(P)$ given the support of the Pauli string $P$. The middle section of brick-wall transfer matrices $T$ (purple) matches the structure of $\mathcal{E}_{U}$. Each transfer matrix $T$ defines a local Markov process, so that the resulting “state” before the final measurement defines a probability distribution over possible supports of the Pauli string. See Eq. (27) for the general form of the transfer matrix. The final layer of measurement gates $T_{M}$ (red) projects the resulting Pauli weight distribution onto ${|0\rangle}+\frac{1}{3}{|1\rangle}$ where-ever a measurement was performed on the quantum state.
@@
Traditionally, the interest in classical shadow tomography has been mostly geared towards experimental applications: predicting linear observables of generic states on quantum devices. However, in this work, we wish to use the general setup of classical shadow tomography as a backdrop to studying operator spreading in dual-unitary circuits. There are two primary reasons we take this approach: one, this scheme is quite general and allows us to adjust the circuit depth $t$, the system size $n$, the type of gate applied, etc.; two, the growth of the Pauli weight, which characterizes the spatial distribution of an operator, is closely related to the measurement channel of the classical shadow scheme; third, operator spreading in this context characterizes the sample complexity of predicting said operator with respect to the particular circuit model via the shadow norm. Hence, in this section, we will introduce the relevant concepts of classical shadow tomography.
@@
Classical shadow tomography was introduced as an efficient protocol for predicting quantum state observables based on few measurements [34]. The scheme works by generating efficient classical representations (or “snapshots”) of the original state, usually stabilizer states, by repeatedly applying a random unitary to the original state and measuring all the qubits afterward. The specific random unitary ensemble that is used (e.g. ${\mathcal{C}\ell}(2)^{\otimes n},{\mathcal{C}\ell}(2^{n}),$ etc.) may be suited for certain types of observables over others.
@@
The protocol consists of two steps: randomized measurement and classical post-processing, as illustrated in Fig. 1(a). Starting with the initial state $\rho$, apply a randomly sampled unitary $U\in\mathcal{E}_{U}$ so that $\rho\rightarrow U\rho U^{\dagger}$. Then perform a projective measurement on the transformed state $U\rho U^{\dagger}$ in the computational basis ($Z$-basis on each qubit independently). The resulting bit-string state ${|b\rangle}$ (labeled by the bit-string $b$ of measurement outcomes), as well as the performed unitary transformation $U$, will be recorded. These two pieces of information define a snapshot state $\hat{\sigma}=U^{\dagger}{|b\rangle}{\langle b|}U$. The randomized measurement protocol can be formulated as a quantum channel $\mathcal{M}$, which maps the initial state $\rho$ to the average snapshot state $\sigma$ by
@@
If the unitary ensemble $\mathcal{E}_{U}$ is local-basis invariant (or equivalently, Pauli-invariant), then the eigen-operators of the measurement channel $\mathcal{M}$ are Pauli strings $P$, with eigenvalue $w_{\mathcal{E}_{\sigma}}$ denoted as the Pauli weight [44, 42, 48, 53]. Furthermore, this Pauli weight is not dependent on the specific Pauli string but on its support. It is named as such because it characterizes the “weight” of an operator on a particular sub-region defined by the support of the Pauli,
@@
is the locally-scrambled shadow norm $\|P\|_{\mathcal{E}_{\sigma}}^{2}$, which characterizes the sample complexity of predicting $P$ using the tomography scheme defined by $\mathcal{E}_{\sigma}$ [42, 48].
@@ II.1.2 Dual-Unitarity in Qubits
We are interested in simulating the evolution of an initial Pauli string $P$ under a brick-wall configuration of random, dual-unitary gates as depicted in Fig. 1. To define the ensemble average and the second moment of random, dual-unitary gates, we must first define what a dual-unitary gate is. Dual-unitary circuits have generated a lot of buzz because they’re local; they’re fast scramblers; they can generate chaotic dynamics; they can be used to form $k$-designs; and, they have simple, tractable expressions for two- and four-point correlation functions, including out-of-time-ordered correlators (OTOCs) [14].
@@
Consider a $n$-qubit system arranged on a lattice with Hilbert space $\mathcal{H}=\otimes_{i}\mathcal{H}_{i}$, where $\dim\mathcal{H}_{i}=2$. Every two-local dual-unitary qubit gate $U$ can be parameterized (up to an overall phase) by four arbitrary single-qubit unitaries $u_{\pm},v_{\pm}$, and a single coefficient $J$, which tunes the coupling in a $XXZ$-type interaction Hamiltonian [56, 57, 58, 59, 60]. This is given below,
@@
We are interested in the average, basis-independent behavior of dual-unitary circuits for prediction in a classical shadow setting. If we choose $u_{\pm},v_{\pm}$ according to a random unitary ensemble, then we are able to define a random, dual-unitary gate ensemble $\mathcal{E}_{U}$.